OpenStudy (anonymous):

Determine any restrictions on the variable. 5y^2/3y-4

OpenStudy (solomonzelman):



OpenStudy (dan815):

you just have to see how there can be a zero in the denominator

OpenStudy (anonymous):

Thank you but can you explain how I work it out? 3y-4=0? Now what?

OpenStudy (dan815):

okay so solve for y

OpenStudy (dan815):

sometimes though it might not be a restriction still so we must check

OpenStudy (anonymous):

Ok I got it. How do I know if it is a restriction or not?

OpenStudy (dan815):

well if it as indeterminante form like if you have 0/0 or inf/inf or 0/inf

OpenStudy (dan815):

in this case we dont have that because the bottom is zero when y = 4/3 but the top which is 5*(4/3)^2 wont be zero so you are fine

OpenStudy (dan815):

means there as an asymtote at y=4/3

OpenStudy (anonymous):

Ok. I have another problem that i am not quite sure how to work out. x+1/x^2-4x?

OpenStudy (dan815):

factor the bottom and see what can make the bottom 0

OpenStudy (dan815):